OpenStudy (anonymous):

Factor completely, then place the answer in the proper location on the grid. a^2 - 25b^2

OpenStudy (campbell_st):

this is the difference of 2 squares the basic form is \[x^2 - y^2 = (x -y)(x + y)\] so just apply this to your question

OpenStudy (anonymous):

i never got to learn this that's why i'm struggling , please help

OpenStudy (campbell_st):

25 is a square number and b^2 is a square so what when multiplied by itself gives 25b^2

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

a^2 - 25b^2 = (a+5b)(a-5b)

OpenStudy (anonymous):

thanks!